# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

### v1

@@ 3. Subset hyperspaces as quotients of function spaces
We will characterize a class of subset hyperspaces as explicit quotients of function spaces. To have reasonable separation properties, we will often choose the function space topologies to be $\tau_{p}$-compatible and likewise choose the hyperspaces topologies to be $\tau_{v}$-compatible.
Definition 3.1 (Indexed subset hyperspaces: Product topology, Limit Vietoris topology, Unordering map).
@@
be their Cartesian product as sets. The product topology $\tau_{p}$ on ${\prod\mathcal{X}}$ is the topology with base sets
@@
Give $X$ the topology $\mathcal{O}(X):=\{T\subset X:T\cap X_{y}\subset X_{y}~{}\textrm{is open}~{}\forall y\in Y\}$, and call it the limit topology on $X$. Give $\mathcal{P}^{\ast}(X):=\mathcal{P}(X)\backslash\{\emptyset\}$ the topology $\tau_{v}$ (and call it the limit Vietoris topology on $\mathcal{P}^{\ast}(X)$) with base sets
$[T_{F}]_{v}=[\{T_{\alpha}:{\alpha}\in F\}]_{v}$ :=$:=\textstyle\{A\in\mathcal{P}^{\ast}(X):A\subset\bigcup_{\alpha}T_{\alpha},~{}A\cap T_{\alpha}\neq\emptyset~{}\forall{\alpha}\},$
for finite collections ${}T_{F}=\{T_{\alpha}:{\alpha}\in F\}\subset\mathcal{O}(X)$. (footnote [^4] The collection :=$\mathcal{B}:=\{[T_{F}]_{v}:F~{}\textrm{finite}\}$ indeed forms a base for a topology on $\mathcal{P}^{\ast}(X)$, because if we let $T_{F}=\{T_{1},...,T_{n}\}$ and $T^{\prime}_{F^{\prime}}=\{T^{\prime}_{1},...,T^{\prime}_{n^{\prime}}\}$, then $[T_{F}]_{v}\cap[T^{\prime}_{F^{\prime}}]_{v}=\big{[}\{(\bigcup_{i}T_{i})\cap T_{1}^{\prime},\cdots,(\bigcup_{i}T_{i})\cap T_{n^{\prime}}^{\prime},T_{1}\cap(\bigcup_{j}T^{\prime}_{j}),\cdots,T_{n}\cap(\bigcup_{j}T^{\prime}_{j})\}\big{]}_{v}=\big{[}\big{\{}(\bigcup T_{F})\cap T^{\prime}_{\alpha^{\prime}}:\alpha^{\prime}\in F^{\prime}\big{\}}\cup\big{\{}T_{\alpha}\cap\bigcup T^{\prime}_{F^{\prime}}:\alpha\in F\big{\}}\big{]}_{v}$, which also lies in $\mathcal{B}$. ).
Let us now define the following items:
@@
    The unordering map $q:{\prod\mathcal{X}}\rightarrow Cl(X),~{}f\mapsto\overline{f(Y)}$.
    Given a $j$-labeled topology $\tau_{j}$ on $\mathcal{G}\subset{\prod\mathcal{X}}$, we denote by $\tau_{jq}$ the topology (called the $\tau_{j}$-quotient topology) on $q(\mathcal{G})\subset Cl(X)$ induced by the restriction $q:(\mathcal{G},\tau_{j})\rightarrow q(\mathcal{G})=(q(\mathcal{G}),\tau_{jq})$ as a quotient map. (footnote555 Recall that $\tau_{jq}$ is the strongest topology $\tau_{jc}$ on $q(\mathcal{G})\subset{\prod\mathcal{X}}$ such that $q:(\mathcal{G},\tau_{j})\rightarrow(q(\mathcal{G}),\tau_{jc})$ is continuous. In particular, if $\tau_{j}:=\tau\cap\mathcal{G}$ (i.e., $\tau_{j}$ is the $\tau$-subspace topology of $\mathcal{G}$ associated with some topology $\tau$ on ${\prod\mathcal{X}}$) then $\tau_{q}\cap q(\mathcal{G})\subset\tau_{jq}=(\tau\cap\mathcal{G})_{q}$ (where $\tau_{q}$ is the $\tau$-quotient topology of $q({\prod\mathcal{X}})$ and $\tau_{q}\cap q(\mathcal{G})$ is the $\tau_{q}$-subspace topology of $q(\mathcal{G})\subset q({\prod\mathcal{X}})$). Therefore, continuity of $q:(\mathcal{G},\tau\cap\mathcal{G})\rightarrow(q(\mathcal{G}),(\tau\cap\mathcal{G})_{q})$ implies continuity of $q:(\mathcal{G},\tau\cap\mathcal{G})\rightarrow(q(\mathcal{G}),\tau_{q}\cap q(\mathcal{G}))$. )
@@
    The set of $Y$-indexed closed subsets of $X$: $Cl_{Y}(\mathcal{X}):=Cl(X)\cap q({\prod\mathcal{X}})\stackrel{{\scriptstyle(\ast)}}{{=}}q({\prod\mathcal{X}})$ (where step ($\ast$) is due to the use of closure in the definition of $q$). Also let
@@
    The set of $Y$-indexed compact subsets of $X$: $K_{Y}(\mathcal{X}):=K(X)\cap q({\prod\mathcal{X}})$. Also let
@@
    The set of $Y$-indexed bounded closed subsets of $X$ (if $X$ is a metric space): $BCl_{Y}(\mathcal{X}):=BCl(X)\cap q({\prod\mathcal{X}})$. Also let
@@
When $\mathcal{X}=\{X\}$, i.e., $X_{y}=X$ for all $y\in Y$, then we will replace $\mathcal{X}$ with $X$ in the $Y$-indexed subset/function spaces above by setting
@@
(3.1) $BCl_{Y}(X):=BCl_{Y}(\mathcal{X}),~{}~{}BCl(Y,X):=BCl(Y,\mathcal{X}).$
Note 3.2.
If $X$ is $T_{1}$ (i.e., singletons of $X$ are closed) then $X\cong F_{1}(X)=\textrm{Singl}(X)$.
Definition 3.3 ($q$-full subset).
In the setup of Definition 3.1, a subset $\mathcal{F}\subset\prod\mathcal{X}$ is $q$-full if $q(\mathcal{F})=q(\prod\mathcal{X})$.
Notice that if $X,Y$ are sets and $\mathcal{F}\subset X^{Y}$ is $q$-full, then we have an injection $X\hookrightarrow\mathcal{F},~{}x\mapsto c_{x}$, where $c_{x}$ is the constant map $Y\rightarrow X,~{}y\mapsto x$. Consequently, if $X$ is a space, then topologies on $\mathcal{F}$ may be seen as extensions/generalizations of the topology of $X$.
Definition 3.4 (rc-topology, rc-space).
In the setup of Definition 3.1, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$ and consider a topology $\tau$ on $\mathcal{F}$. A net $f_{\alpha}\in(\mathcal{F},\tau)$ is compactly-ranged (hence a cr-net) if there exists a compact set $K\subset X=\bigcup_{y\in Y}X_{y}$ and a tail $T_{\beta}:=\{f_{\alpha}:\alpha\geq\beta\}$ of $f_{\alpha}$ such that $\bigcup T_{\beta}:=\bigcup_{\alpha\geq\beta}q(f_{\alpha})\subset K$.
@@
Definition 3.6 (wrc-topology, wrc-space).
In the setup of Definition 3.1, let $\tau^{\prime}$ be a topology on $Cl_{Y}(\mathcal{X})$, and fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A topology $\tau$ on $\mathcal{F}$ is $\tau^{\prime}$-weakly range-compact (hence a $\tau^{\prime}$-wrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau)$ a $\tau^{\prime}$-wrc-space) if for any $\tau^{\prime}$-convergent net $C_{\alpha}\in(K_{Y}(\mathcal{X}),\tau^{\prime})$ such that $\bigcup_{\alpha\geq\beta}C_{\alpha}\subset K$ for a compact set $K\subset X$ and some index $\beta$, every net $g_{\alpha}\in q^{-1}(C_{\alpha})$ has a $\tau$-convergent subnet in $(\mathcal{F},\tau)$. (footnote666 Observe that the rc-property is stronger than (i.e., implies) the wrc-property. That is, every rc-topology (such as $\tau_{p}$) is a wrc-topology. ).
@@
Definition 3.7 (swrc-topology, swrc-space, Standard topology of the indexed subset hyperspaces).
In the setup of Definition 3.1, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A wrc-topology $\tau_{\pi}$ on $\mathcal{F}$ is a standard wrc-topology (hence a swrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau_{\pi})$ a swrc-space) if $\tau_{v}\subset\tau_{\pi q}$ in $K_{Y}(\mathcal{X})$ (where $\tau_{\pi q}$ denotes the $\tau_{\pi}$-quotient topology on $Cl_{Y}(\mathcal{X})$), i.e., if with
@@
Question 3.8.
Let $X$ be a space, $Y$ a set (and let $\mathcal{X}:=\{X\}$), and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. How do we explicitly specify the preferred topology $\tau_{\pi}$ of Definition 3.7? Theorem 4.4 specifies $\tau_{\pi}$ for some special cases where $\mathcal{F}$ is sufficiently well-behaved.
Definition 3.9 (Set-open topology, Compact-open topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{S}\subset\mathcal{P}(Y)$ a family of subsets of $Y$. The $\mathcal{S}$-open topology $\tau_{\mathcal{S}}$ on $X^{Y}$ is the topology with subbase given by the sets
$[S,O]_{\mathcal{S}}:=\{f\in X^{Y}:f(S)\subset O\},~{}~{}~{}~{}\textrm{for sets$S\in\mathcal{S}$ and open sets $O\subset X$}.$
@@
    Based on the setup in Definition 3.7, if $X,Y$ are spaces and $\mathcal{F}\subset X^{Y}$ is $q$-full, we have the quotient map
@@
    In this special case (where both $X$ and $Y$ are spaces), candidates for $\tau_{\pi}$ in Definition 3.7 are $q^{-1}(\tau_{v})$ and $\tau_{co}$, and it might therefore be reasonable/sufficient (regarding Question 3.8) to search for a $\tau_{p}$-compatible $\tau_{\pi}$ in the range (footnote777 Consider the map $q:(X^{Y},\tau_{1},\tau_{2})\rightarrow(Cl_{Y}(X),\tau_{1q},\tau_{2q})$. If $\tau_{1}\subset\tau_{2}$ then $B\in\tau_{1q}$ $\iff$ $q^{-1}(B)\in\tau_{1}\subset\tau_{2}$, $\Rightarrow$ $q^{-1}(B)\in\tau_{2}$, $\iff$ $B\in\tau_{2q}$ (i.e., $\tau_{1}\subset\tau_{2}$Rightarrow $\tau_{1q}\subset\tau_{2q}$). )
@@
    (in accordance with   $\tau_{v}\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}\subset\tau_{\pi q}\subset\tau_{co\!\!\!~{}q}$   due to Theorem 3.17 and [bartsch2014, Cor. 2.3]), where $\tau_{s}$ is the topology of symmetric convergence (Definition 3.12).
@@
Definition 3.11 (Vietoris topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. For any open set $O\subset X$ and any finite collection of open sets $O_{F}=\{O_{\alpha}:{\alpha}\in F\}$ in $X$, let
$[O_{F}]_{v}$ :=$\textstyle:=\{C\in Cl_{Y}(X):C\subset\bigcup_{{\alpha}\in F}O_{\alpha},~{}C\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\}$
=$\textstyle=\left\{q(f):f\in\mathcal{F},~{}q(f)\subset\bigcup_{{\alpha}\in F}O_{\alpha},~{}q(f)\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\right\},$
$O^{+}$ :=[$\textstyle:=[\{O\}]_{v}=\{C\in Cl_{Y}(X):C\subset O\}=\left\{q(f):f\in\mathcal{F},~{}q(f)\subset O\right\},$
$O^{-}$ :=[$\textstyle:=[\{X,O\}]_{v}=[\{O^{c}\}]_{v}{}^{c}=\{C\in Cl_{Y}(X):C\cap O\neq\emptyset\}=\left\{q(f):f\in\mathcal{F},~{}q(f)\cap O\neq\emptyset\right\}$
=$\textstyle=\left\{q(f):f\in\mathcal{F},~{}f(Y)\cap O\neq\emptyset\right\}.$
The Vietoris topology $\tau_{v}$ (also see Definition 3.1) of $Cl_{Y}(X)$ is the topology with base
@@
$S\mathcal{B}_{v}^{+}=\big{\{}O^{+}:\textrm{$O\subset X$ open}\big{\}}~{}~{}\big{(}\textrm{resp},~{}~{}S\mathcal{B}_{v}^{-}=\big{\{}O^{-}:\textrm{$O\subset X$ open}\big{\}}\big{)}.$
@@
Definition 3.12 (Topology of symmetric convergence).
@@
$[O]_{s}:=\{f\in X^{Y}:q(f)\subset O\},~{}~{}~{}~{}\textrm{for open sets}~{}O\subset X.$
Definition 3.13 (Finitely $q$-stable set of functions).
@@
=$\textstyle\stackrel{{\scriptstyle(\ast)}}{{=}}\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~{}f(Y)\cap O_{y}^{i}\neq\emptyset,\forall y\in F_{i}\right\}$
@@
=$\textstyle=\left\{g\in\mathcal{F}:g(Y)\cap O\neq\emptyset\right\}=\left\{g\in\mathcal{F}:\exists y\in Y,g(y)\in O\right\}=\bigcup_{y\in Y}\left\{g\in\mathcal{F}:g(y)\in O\right\}$
@@
(ii) By the above relation, unless $Y$ is a singleton, we have the proper containment ${}q^{-1}(\tau_{v}^{-})\subsetneq\tau_{p}$ (which holds because, by construction, both $q^{-1}(O^{-})$ and $q^{-1}(O^{+})$ cannot distinguish between the points of $Y$). ∎
Remark 3.15.
@@
    =$\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}=\left\{g\in\mathcal{F}:g(Y)\subset O\right\}\cap\left\{g\in\mathcal{F}:\partial g(Y)\subset O\right\}$
@@
    $\tau_{gq}=\tau_{v}^{+}$ in $Cl_{Y}(X)$ (i.e., the map $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is a quotient map),
@@
    If $X$ is $T_{1}$, then $\tau_{s}\subset\tau_{p}$ $\iff$ $Y$ is finite (by Remark 3.15(2)).
@@
(i) Proving $\tau_{gq}\subset\tau_{v}^{+}$: A set $\mathcal{A}\subset Cl_{Y}(X)$ is $\tau_{gq}$-open iff $q^{-1}(\mathcal{A})=\left\{f\in\mathcal{F}:q(f)\in\mathcal{A}\right\}$ is $\tau_{s}$-open, i.e., iff there exist a collection of open sets $\{O_{i}\subset X:i\in I\}$ such that
@@
Proving $\tau_{gq}\supset\tau_{v}^{+}$ (i.e., $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is continuous): As in Remark 3.15(2),
@@
=$\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}\in\tau_{s}.$

### v2

@@ 3. Subset hyperspaces as quotients of function spaces
We will characterize a class of subset hyperspaces as explicit quotients of function spaces. To have reasonable separation properties, we aim to choose the function space topologies to be $\tau_{p}$-compatible and likewise choose the hyperspaces topologies to be $\tau_{v}$-compatible.
Definition 3 (Indexed subset hyperspaces: Limit Vietoris topology, Unordering map).
@@
be their Cartesian product as sets. Recall that the product topology $\tau_{p}$ on $\prod\mathcal{X}$} has base sets
@@
[…]
$[T_{F}]_{v}=[\{T_{\alpha}:{\alpha}\in F\}]_{v}$ $:=\textstyle\{A\in\mathcal{P}^{\ast}(X):A\subset\bigcup_{\alpha}T_{\alpha},~A\cap T_{\alpha}\neq\emptyset~\forall{\alpha}\in F\},$
for finite collections $T_{F}=\{T_{\alpha}:{\alpha}\in F\}\subset\mathcal{O}(X)$, where $F$ is an arbitrary finite set. (footnote [^3] The collection $\mathcal{B}:=\{[T_{F}]_{v}:F~\textrm{finite}\}$ indeed forms a base for a topology on $\mathcal{P}^{\ast}(X)$, because if we let $T_{F}=\{T_{1},...,T_{n}\}$ and $T^{\prime}_{F^{\prime}}=\{T^{\prime}_{1},...,T^{\prime}_{n^{\prime}}\}$, then $T_{F}]_{v}\cap[T^{\prime}_{F^{\prime}}]_{v}=\big[\{(\bigcup_{i}T_{i})\cap T_{1}^{\prime},\cdots,(\bigcup_{i}T_{i})\cap T_{n^{\prime}}^{\prime},T_{1}\cap(\bigcup_{j}T^{\prime}_{j}),\cdots,T_{n}\cap(\bigcup_{j}T^{\prime}_{j})\}\big]_{v}=\big[\big\{(\bigcup T_{F})\cap T^{\prime}_{\alpha^{\prime}}:\alpha^{\prime}\in F^{\prime}\big\}\cup\big\{T_{\alpha}\cap\bigcup T^{\prime}_{F^{\prime}}:\alpha\in F\big\}\big]_{v}$, which also lies in $\mathcal{B}$. ).
Let us now define the following items (where items (1)-(4) have already appeared in the introduction under a less general setup):
@@
    The unordering map $q:{\prod\mathcal{X}}\rightarrow Cl(X),~f\mapsto cl_{X}(f(Y))$.
    Given a $j$-labeled topology $\tau_{j}$ on $\mathcal{F}\subset{\prod\mathcal{X}}$, we denote by $\tau_{jq}$ the topology (called the $\tau_{j}$-quotient topology) on $q(\mathcal{F})\subset Cl(X)$ induced by the restriction $q:(\mathcal{F},\tau_{j})\rightarrow q(\mathcal{F})=(q(\mathcal{F}),\tau_{jq})$ as a quotient map. (footnote [^4] Recall that $\tau_{jq}$ is the strongest topology $\tau_{jc}$ on $q(\mathcal{F})\subset{\prod\mathcal{X}}$ such that $q:(\mathcal{F},\tau_{j})\rightarrow(q(\mathcal{F}),\tau_{jc})$ is continuous. In particular, if $\tau_{j}:=\tau\cap\mathcal{F}$ (i.e., $\tau_{j}$ is the $\tau$-subspace topology of $\mathcal{F}$ associated with some topology $\tau$ on $\prod\mathcal{X}$}) then $\tau_{q}\cap q(\mathcal{F})\subset\tau_{jq}=(\tau\cap\mathcal{F})_{q}$ (where $\tau_{q}$ is the $\tau$-quotient topology of $q({\prod\mathcal{X}})$ and $\tau_{q}\cap q(\mathcal{F})$ is the $\tau_{q}$-subspace topology of $q(\mathcal{F})\subset q({\prod\mathcal{X}})$). Therefore, continuity of $q:(\mathcal{F},\tau\cap\mathcal{F})\rightarrow(q(\mathcal{F}),(\tau\cap\mathcal{F})_{q})$ implies continuity of $q:(\mathcal{F},\tau\cap\mathcal{F})\rightarrow(q(\mathcal{F}),\tau_{q}\cap q(\mathcal{F}))$. )
@@
    The set of $Y$-indexed closed subsets of $X$: $Cl_{Y}(\mathcal{X}):=Cl(X)\cap q({\prod\mathcal{X}})\stackrel{{\scriptstyle(\ast)}}{{=}}q({\prod\mathcal{X}})$ (where step ($\ast$) is due to the use of closure in the definition of $q$). Also, let
@@
    The set of $Y$-indexed compact subsets of $X$: $K_{Y}(\mathcal{X}):=K(X)\cap q({\prod\mathcal{X}})$. Also, let
@@
    The set of $Y$-indexed bounded closed subsets of $X$ (if $X$ is a metric space): $BCl_{Y}(\mathcal{X}):=BCl(X)\cap q({\prod\mathcal{X}})$. Also, let
@@
When $\mathcal{X}=\{X\}$, i.e., $X_{y}=X$ for all $y\in Y$, we will replace $\mathcal{X}$ with $X$ in the $Y$-indexed subset/function spaces above by setting
@@
$BCl_{Y}(X):=BCl_{Y}(\mathcal{X}),~~BCl(Y,X):=BCl(Y,\mathcal{X}).$
Note 1.
If $X$ is $T_{1}$ (i.e., singletons of $X$ are closed in $X$) then $X\cong F_{1}(X)=\textrm{Singlt}(X)$.
Definition 4 ($q$-full subset).
In the setup of Definition 3, a subset $\mathcal{F}\subset\prod\mathcal{X}$ is $q$-full if
$q(\mathcal{F})=q(\prod\mathcal{X}).$
Notice that if $X,Y$ are sets and $\mathcal{F}\subset X^{Y}$ is $q$-full, then we have an injection $X\hookrightarrow\mathcal{F},~x\mapsto c_{x}$, where $c_{x}$ is the constant map $Y\rightarrow X,~y\mapsto x$. Consequently, if $X$ is a space, then topologies on $\mathcal{F}$ (just like hypertopologies of a $T_{1}$ space $X$) may be seen as extensions/generalizations of the topology of $X$.
Definition 5 (rc-topology, rc-space).
In the setup of Definition 3, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$ and consider a topology $\tau$ on $\mathcal{F}$. A net $f_{\alpha}\in(\mathcal{F},\tau)$ is compactly-ranged (hence a cr-net) if there exists a compact set $K\subset X=\bigcup_{y\in Y}X_{y}$ and a tail $T_{\beta}:=\{f_{\alpha}:\alpha\geq\beta\}$ of $f_{\alpha}$ such that $\bigcup q(T_{\beta}):=\bigcup_{\alpha\geq\beta}q(f_{\alpha})\subset K$.
@@
Definition 6 (wrc-topology, wrc-space).
In the setup of Definition 3, let $\tau^{\prime}$ be a topology on $Cl_{Y}(\mathcal{X})$, and fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A topology $\tau$ on $\mathcal{F}$ is $\tau^{\prime}$-weakly range-compact (hence a $\tau^{\prime}$-wrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau)$ a $\tau^{\prime}$-wrc-space) if for any $\tau^{\prime}$-convergent net $C_{\alpha}\in(K_{Y}(\mathcal{X}),\tau^{\prime})$ such that $\bigcup_{\alpha\geq\beta}C_{\alpha}\subset K$ for a compact set $K\subset X$ and some index $\beta$, every net $g_{\alpha}\in q^{-1}(C_{\alpha})$, $\alpha\geq\beta$, has a $\tau$-convergent subnet in $(\mathcal{F},\tau)$. (footnote [^5] Observe that the rc-property is stronger than (i.e., implies) the wrc-property. That is, every rc-topology (such as $\tau_{p}$) is a wrc-topology. ).
@@
Definition 7 (swrc-topology, swrc-space, Standard topology of the indexed subset hyperspaces).
In the setup of Definition 3, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A wrc-topology $\tau_{\pi}$ on $\mathcal{F}$ is a standard wrc-topology (hence a swrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau_{\pi})$ a swrc-space) if $\tau_{v}\subset\tau_{\pi q}$ in $K_{Y}(\mathcal{X})$ (where $\tau_{\pi q}$ denotes the $\tau_{\pi}$-quotient topology on $Cl_{Y}(\mathcal{X})$), i.e., if with
@@
Question 1.
Let $X$ be a space, $Y$ a set (and let $\mathcal{X}:=\{X\}$), and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. How do we explicitly specify the preferred topology $\tau_{\pi}$ of Definition 7? Theorem 4.2 specifies $\tau_{\pi}$ for some special cases where $\mathcal{F}$ is sufficiently well-behaved.
Definition 8 (Set-open topology, Compact-open topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{S}\subset\mathcal{P}(Y)$ a family of subsets of $Y$. The $\mathcal{S}$-open topology $\tau_{\mathcal{S}}$ on $X^{Y}$ is the topology with a subbase given by the sets
$S,O]_{\mathcal{S}}:=\{f\in X^{Y}:f(S)\subset O\},~~~~\textrm{for sets $S\in\mathcal{S}$ and open sets $O\subset X$}.$
@@
    Based on the setup in Definition 7, if $X,Y$ are spaces and $\tau_{\pi}$ is a swrc-topology on a $q$-full $\mathcal{F}\subset X^{Y}$, then we have the quotient map
@@
    In this special case (where both $X$ and $Y$ are spaces), candidates for $\tau_{\pi}$ in Definition 7 are $q^{-1}(\tau_{v})$ and $\tau_{co}$, and it might therefore be reasonable/sufficient (regarding Question 1) to search for a $\tau_{p}$-compatible $\tau_{\pi}$ in the range (footnote [^6] Consider the map $q:(X^{Y},\tau_{1},\tau_{2})\rightarrow(Cl_{Y}(X),\tau_{1q},\tau_{2q})$. If $\tau_{1}\subset\tau_{2}$, then $B\in\tau_{1q}$ $\iff$ $q^{-1}(B)\in\tau_{1}\subset\tau_{2}$, $\Rightarrow$ $q^{-1}(B)\in\tau_{2}$, $\iff$ $B\in\tau_{2q}$ (i.e., $\tau_{1}\subset\tau_{2}$ $\Rightarrow$ $\tau_{1q}\subset\tau_{2q}$). )
@@
    (in accordance with   $\tau_{v}\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}\subset\tau_{\pi q}\subset\tau_{co\!\!\!~q}$   due to Theorem 3.3 and [bartsch2014, Corollary 2.3]), where $\tau_{s}$ is the topology of symmetric convergence (Definition 10).
@@
We will now discuss the Vietoris topology (a more general version of which we have already introduced in Definition 3) in sufficient detail for our subsequent discussion. Using convergence of nets, a more intuitive interpretation of the Vietoris topology is given later in Definition 13.
Definition 9 (Vietoris topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. For any open set $O\subset X$ and any finite collection of open subsets $O_{F}=\{O_{\alpha}:{\alpha}\in F\}$ of $X$, let
$[O_{F}]_{v}$ $\textstyle:=\{C\in Cl_{Y}(X):C\subset\bigcup_{{\alpha}\in F}O_{\alpha},~C\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\}$
$\textstyle=\left\{q(f):f\in\mathcal{F},~q(f)\subset\bigcup_{{\alpha}\in F}O_{\alpha},~q(f)\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\right\},$
$O^{+}$ $\textstyle:=[\{O\}]_{v}=\{C\in Cl_{Y}(X):C\subset O\}=\left\{q(f):f\in\mathcal{F},~q(f)\subset O\right\},$
$O^{-}$ $\textstyle:=[\{X,O\}]_{v}=[\{O^{c}\}]_{v}{}^{c}=\{C\in Cl_{Y}(X):C\cap O\neq\emptyset\}=\left\{q(f):f\in\mathcal{F},~q(f)\cap O\neq\emptyset\right\}$
$\textstyle=\left\{q(f):f\in\mathcal{F},~f(Y)\cap O\neq\emptyset\right\}.$
The Vietoris topology $\tau_{v}$ (also see Definition 3) of $Cl_{Y}(X)$ is the topology with base
@@
$S\mathcal{B}_{v}^{+}=\big\{O^{+}:\textrm{$O\subset X$ open}\big\}~~\big(\textrm{resp.},~~S\mathcal{B}_{v}^{-}=\big\{O^{-}:\textrm{$O\subset X$ open}\big\}\big).$
@@
Definition 10 (Topology of symmetric convergence).
@@
$O]_{s}:=\{f\in X^{Y}:q(f)\subset O\},~~~~\textrm{for open sets}~O\subset X.$
Definition 11 (Finitely $q$-stable set of functions).
@@
$\textstyle\stackrel{{\scriptstyle(\ast)}}{{=}}\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~f(Y)\cap O_{y}^{i}\neq\emptyset,\forall y\in F_{i}\right\}$
@@
$\textstyle=\left\{g\in\mathcal{F}:g(Y)\cap O\neq\emptyset\right\}=\left\{g\in\mathcal{F}:\exists y\in Y,g(y)\in O\right\}=\bigcup_{y\in Y}\left\{g\in\mathcal{F}:g(y)\in O\right\}$
@@
(ii) This follows from the observation that the subbase elements $\{[y,O]_{p}:y\in Y\}$ of $\tau_{p}$ can distinguish the points of $Y$, meanwhile, by construction, neither the elements of $q^{-1}(\tau_{v}^{-})$ nor those of $q^{-1}(\tau_{v}^{+})$ can distinguish the points of $Y$. Indeed, if $|Y|\geq 2$, pick $y_{1},y_{2}\in Y$ such that $y_{1}\neq y_{2}$ and $y_{1},O]_{p}\neq[y_{2},O]_{p}$. Then each $y_{i},O]_{p}$ (an element of $\tau_{p}$) depends asymmetrically on $y_{1}$ and $y_{2}$. But every member of $q^{-1}(\tau^{-}_{v})$ depends symmetrically on $y_{1}$ and $y_{2}$ (as the expression for $q^{-1}(O^{-})$ above shows), and so $q^{-1}(\tau^{-}_{v})$ does not contain $y_{i},O]_{p}$. That is, if $|Y|\geq 2$, then $\tau_{p}\backslash q^{-1}(\tau_{v}^{-})\neq\emptyset$ (provided $X$ is a sufficiently nontrivial space). ∎
Remark.
@@
    $\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}=\left\{g\in\mathcal{F}:g(Y)\subset O\right\}\cap\left\{g\in\mathcal{F}:\partial g(Y)\subset O\right\}$
@@
    $\tau_{sq}=\tau_{v}^{+}$ in $Cl_{Y}(X)$ (i.e., the map $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is a quotient map),
@@
    If $X$ is $T_{1}$, then $\tau_{s}\subset\tau_{p}$ $\iff$ $Y$ is finite (by Lemma 3.1’s Remark (2)).
@@
(i) Proving $\tau_{sq}\subset\tau_{v}^{+}$: A set $\mathcal{A}\subset Cl_{Y}(X)$ is $\tau_{sq}$-open iff $q^{-1}(\mathcal{A})=\left\{f\in\mathcal{F}:q(f)\in\mathcal{A}\right\}$ is $\tau_{s}$-open, i.e., iff there exist a collection of open sets $\{O_{i}\subset X:i\in I\}$ such that
@@
Proving $\tau_{sq}\supset\tau_{v}^{+}$ (i.e., $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is continuous): As in Lemma 3.1’s Remark (2),
@@
$\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}\in\tau_{s}.$